OpenStudy (anonymous):

suppose S is between R and T. Use the segment addition postulate to solve for each variable. RS= 7y-4 , ST= y+5 , RT= 28

OpenStudy (anonymous):

Look at this as three seperate lines at first to visualize the problem. RS, ST, and RT. You know that the entire segment is RT and S is just somewhere in there somewhere. r--------------s----------------------... (my attempt at drawing) Based on this model you can see that RS + ST make up the entire segment RT so set RS + ST = RT (7y - 4) + (y+5) = 28 combine terms 7y and y and -4 and 5 8y + 1 = 28 subtract 1 from each side 8y = 27 divide by 8 y = 27/8

OpenStudy (anonymous):

@Kbdestroy